OpenStudy (anonymous):

Holly wants to save money for an emergency. Holly invests $500 in an account that pays an interest rate of 4%. How many years will it take for the account to reach 12,000? Round answer to nearest hundredth.

OpenStudy (anonymous):

Interest= principle x rate x time (in years) 11,500=(500)(.04)t 11,500=20t 575=t 575 years

OpenStudy (wolf1728):

Or a neater solution is to use the formula locates here: http://www.1728.org/compint2.htm Years = {log(total) -log(Principal)} ÷ log(1 + rate) Years = log (12,000) -log(500) ÷ log (1.04) Years = (4.079181246 - 2.6989700043) ÷ 0.0170333393 Years = 1.3802112417 ÷ 0.0170333393 Years = 81.029985812 Here's a link to a calculator http://www.1728.org/compint.htm Enter the numbers and the answer is 81.03